OpenStudy (anonymous):

What is the equation of a line, in general form, that passes through point (1, -2) and has a slope of 1/3 <-----thats a fraction

OpenStudy (alekos):

you got any ideas?

OpenStudy (anonymous):

@alekos Not a clue, I was doing it all wrong and now im stuck

OpenStudy (alekos):

you need to express in the form y = mx + b where m is the slope and b is the y intercept does that help?

OpenStudy (anonymous):

@alekos these are my choices... Im thinking B A. x - y + 2 = 0 B. x + y - 2 = 0 C. x - y - 2 = 0

OpenStudy (alekos):

my mistake, has to be in general form Ax + By + C = 0 slope would be - A/B in this case you're given 1/3 so A would be -1 and B would be 3 so none of the choices are valid